I cannot get this please help
solve using the substitution method
4m+n=27
m-3n=23

Answer by mathsmiles(68)

(I think you're the same person I just helped with another substitution.)
What we want is to get one of the variables alone on one side of the equation. Looking at the two equations, my choice is getting M by itself in the 2nd equation.
m-3n=23
Add 3n to both sides and wallah!
+3n +3n
m = 23 + 3n

Now, "substitute" this for m in the first equation (aha you say!)
4m + n = 27
4(23+3n) + n = 27
92 + 12n + n = 27
92 + 13n = 27
subtract 92 from both sides:
-92 -92
13n = -65
Divide both sides by 13:
n = -5

SOOOO,
m = 23 + 3n
m = 23 + 3(-5)
m = 23 - 15
m = 8

Checking:
4(8) + -5 = 27
32 - 5 = 27
27 = 27 Correct
8-3(-5) = 23
8 + 15 = 23
23 = 23 Correct!

I wish I had your textbook to look up elimination method and I could help you with that. It's not the term I used and I'm afraid I'll misguide you. If you could look that up and add it to one of your "elimination" method questions, I'll stay online a little longer.
Smile! You're going to get this! You now know "substitution". :-)